Problem 12 Break-even with multiple products Boston Ballet is located in Boston. The company is housed in the Opera House in Boston downtown. The company features five different ballets per year: The Dream, Peter Pan, The Nutcracker, Sleeping Beauty, and Swan Lake. The tentative scheduled for each ballet is as following: heam Peter Pan Nutcracker Sleeping Beauty Swan Lake 20 10 To produce each ballet, there are costs for costumes, props, rehearsals, royalties, guest artist fees, choreography, salaries of production staff, music, and wardrobe. These costs are fixed for a particular ballet regardless of the number of performances. These direct fixed costs are-given below for each ballet. Dream Peter Pan Nutcracker Sleeping Beauty 345,000 275,500 70,500 Swan Lake Other fixed costsfor the whole company are as follows: Advertising 80,000 15,000 222,000 Insurance Administrative Salaries Office rental, phone, etc Total 401,000 For each performance of each ballet, the following costs are incurred: Symphony 3,800 Opera House rental700 Dancers payroll Total 8,500
10 The Opera House in which the ballet is presented has 1,354 scats, which are classified as A, B, and C. The best viewing ranges from A seats to C seats. Information concerning the different types of seats is given below. SeatsBSeats Seats 114 S35 756 $25 984 $15 Quantity Price Estimated percentage sold for each performance Nutcracker Other 4 ballets 100% 100% 100% 80% 100% 75% Questions 1. Prepare a variable costing income statement for each ballet. 2. Calculate the number of performances to reach the break-even for each ballet. 3. Calculate the number of performances of each ballet required for the company to break even, holding the mix of performances constant. Suppose that it is possible to offer an afternoon performance of the popular Nutcracker. Seats would sell for $5 less than in the evening, but the rental of the Opera House would be $200 less. Numbers of seats sold for each type would be 80%. Irs afternoon performances are 4. added, what will be the effect on the company profit?
Answer #1
INCOME STATEMENT
DREAM:
Sales Revenue $150,900 ((114*35)+(0.8*756*25)+(0.75*984*15))*5
Variable Costs $42,500 (5*8500)
Contribution Margin $108,400
Fixed Costs $275,500
Operating income ($167,100)
PETER PAN
Sales Revenue $150,900 ((114*35)+(0.8*756*25)+(0.75*984*15))*5
Variable Costs $42,500 (5*8500)
Contribution Margin $108,400
Fixed Costs $145,500
Operating income ($37,100)
NUTCRACKER
Sales Revenue $753,000 ((114*35)+(756*25)+(984*15))*20
Variable Costs $170,000 (20*8500)
Contribution Margin $583,000
Fixed Costs $70,500
Operating income $512,500
SLEEPING BEAUTY
Sales Revenue $301,800 ((114*35)+(0.8*756*25)+(0.75*984*15))*10
Variable Costs $85,000 (10*8500)
Contribution Margin $216,800
Fixed Costs $345,000
Operating income ($128,200)
SWAN LAKE
Sales Revenue $150,900 ((114*35)+(0.8*756*25)+(0.75*984*15))*5
Variable Costs $42,500 (5*8500)
Contribution Margin $108,400
Fixed Costs $155,500
Operating income ($47,100)
Number of performance to reach Break Even for each ballet
DREAM:
A Contribution Margin $108,400
B Contribution per performance $21,680 (108400/5)
C Fixed Costs $275,500
D=C/B Break even performance                   13 Rounded to next higher round number
PETER PAN
A Contribution Margin $108,400
B Contribution per performance $21,680 (108400/5)
C Fixed Costs $145,500
D=C/B Break even performance                     7 Rounded to next higher round number
NUTCRACKER
A Contribution Margin $583,000
B Contribution per performance $29,150 (583000/20)
C Fixed Costs $70,500
D=C/B Break even performance                     3 Rounded to next higher round number
SLEEPING BEAUTY
A Contribution Margin $216,800
B Contribution per performance $21,680 (216800/10)
C Fixed Costs $345,000
D=C/B Break even performance                   16 Rounded to next higher round number
SWAN LAKE
A Contribution Margin $108,400
B Contribution per performance $21,680 (108400/5)
C Fixed Costs $155,500
D=C/B Break even performance                     8 Rounded to next higher round number
MIX Of Performance Proportion Dream Peter pan Nutcracker Sleeping Beauty Swan lake Total 10 Unit Weighted Weight in the Mix Dream Peter pan Nutcracker Sleeping Beauty Swan lake eightContribution Contributiorn 0.10 0.20 0.40 0.20 0.10 $21,680 $2,168 $21,680 $4,336 29,150$11,660 $21,680 $4,336 $21,680 $2,168 $24,668 Tota Total fixed Costs Fixed cost for the company as whole $401,000 Total fixed Costs Breakeven Quantity for the mix $992,000 (275500+145500+70500+345000+145500) 56.47 (1393000/24668) Break Even performance Dream Peter pan Nutcracker Sleeping Beauty Swan lake 6 (56.47*0.1) 12 (56.47*0.2) 23 (56.47*0.4) 12 (56.47*0.2) 6 (56.47*0.1)